# A response matrix determined with a coincidence-based acquisition for correction of charge sharing spectral distortions in energy-resolved photon counting detectors

### v1

@@ 1 Introduction
In this paper a statistical post-processing method for the characterization and correction of charge sharing distortions of PCD X-rays spectra is proposed. The method is based on the collection of calibration data with a reference polychromatic spectrum and a uniform illumination of the detector to determine the elements of a spectral response matrix. At this stage it is required that the number of coincidences between the signal of one pixel and the analog sum of its neighbouring pixels are counted for all the possible combinations of energy bins. Assuming an equilised behaviour of all the PCD pixels, the calibration requires a simple coincidence circuit involving only a limited number of pixels. The calibration allows to define the probabilities that the initial energy is split between one pixel and its neighbours. These probabilities are used to define the elements of a matrix, named “coincidence-based response matrix” (CBRM), which inversion allows to retrieve the input energy distribution in the absence of charge sharing. The CBRM is independent of the input spectrum and can therefore be used to correct CS distortions in subsequent acquisitions with different spectra. The correction requires only the integral number of counts collected for different energy bins by each pixel. Therefore, once the response matrix is determined, the image acquisitions can be performed with conventional multi-comparator readout electronics, without introducing additional dead-times. Compared to other semi-empirical post-processing algorithms, the method is based only on data inputs, is independent of models or simulations, is simple and of general applicability.
@@ 2.2.1 Charge sharing model
    An event in the reference pixel with deposited energy $i$ is shared with one of the neighbouring pixels, reducing the number of counts at bin $i$. This CS-OUT transition has probability $p^{(out)}_{j,i-j}$, where $j$ is the residual energy in the reference pixel and iji-j is the energy released in the nearby pixels.
@@ 2.2.3 Reduction of the CBRM matrix
Eq.16 defines the elements of a new $L\times L$ response matrix $A^{(r)}$, adapted to the wider energy bins of the multi-comparator system with a limited number of energy thresholds. The inversion of the reduced response matrix, or the iterative solution of Eq.16, allows to retrieve the input events $m_{i}$.
@@ 2.3 Simulation of a CdTe detector
In order to validate the method described in the previous section, the response of a pixelated CdTe detector of 1 mm thickness and segmented in 9x9 pixels was simulated with the Geant4 toolkit (Agostinelli et al., 2003) version 11.1.3. Different pixel sizes were considered, between 100 and 300 µm. A proper number of X-ray photons were generated with a direction perpendicular to the detector, and with an energy randomly selected from a polychromatic X-ray spectrum produced with the Spektr code (Punnoose et al., 2016). The photon source was placed at a fixed distance from the detector with transversal coordinates uniformly distributed to illuminate the whole detector.
@@ 3.1 Simulation studies
Figure 5: MAPE values from the comparison of the CBRM correction (left) and the ACS correction (right) with the reference spectrum of an ideal detector, as a function of the pixel size and the noise discrimination threshold. The input spectrum from a 120 kVp W tube was attenuated by 120 mm of H2O and 0.2 mm of Iodine, and a random noise of 1 keV rms was added to the output of each pixel. The results were obtained from Geant4 simulations.
The effectiveness of the CBRM correction depends also on the number of energy bins used for the definition of the response matrix. In general, a high number of energy bins is desirable for an accurate determination of the matrix; this can be obtained by changing the thresholds of the two discriminators of Fig.1(right) in small steps. However, the typical number of comparators and the number of energy bins available in standard pulse-height acquisition systems is limited. The procedure described in Sect.2.2.3 can be followed to reduce the response matrix to match the number of energy bins of a standard acquisition.
@@
3.2 Study of data from Timepix4 acquisitions
The analysis of the data collected at the SYRMEP beamline was performed using energy bins of 0.5 keV width. The response matrix was determined by merging all the data collected at 18 different energies, to define the transition probabilities over a wide energy range. The CBRM matrix was therefore applied to recover the spectrum of each single acquisition, and the results were compared with the spectra obtained from a clustering algorithm based on 3x3 pixel blocks (Sect.2.5).

### v2

@@ 1 Introduction
In this paper a statistical post-processing method for the characterization and correction of charge sharing distortions of PCD X-ray spectra is proposed. The method is based on the collection of calibration data with a reference polychromatic spectrum and a uniform illumination of the detector to determine the elements of a spectral response matrix. At this stage it is required that the number of coincidences between the signal of one pixel and the analog sum of its neighbouring pixels are counted for all the possible combinations of energy bins. Assuming an equilised behaviour of all the PCD pixels, the calibration requires a simple coincidence circuit involving only a limited number of pixels. The calibration allows to define the probabilities that the initial energy is split between one pixel and its neighbours. These probabilities are used to define the elements of a matrix, named “coincidence-based response matrix” (CBRM), which inversion allows to retrieve the input energy distribution in the absence of charge sharing. The CBRM is independent of the input spectrum and can therefore be used to correct CS distortions in subsequent acquisitions with different spectra. The correction requires only the integral number of counts collected for different energy bins by each pixel. Therefore, once the response matrix is determined, the image acquisitions can be performed with conventional multi-comparator readout electronics, without introducing additional dead-times. Compared to other semi-empirical post-processing algorithms, the method is based only on data inputs, is independent of models or simulations, is simple and of general applicability.
@@ 2.2.1 Charge sharing model
    An event in the reference pixel with deposited energy $i$ is shared with the neighbouring pixels, reducing the number of counts at bin $i$. This CS-OUT transition has probability $p^{(out)}_{j,i-j}$, where $j$ is the residual energy in the reference pixel and iji-j is the energy released in the nearby pixels.
@@ 2.2.3 Reduction of the CBRM matrix
Eq.16 defines the elements of a new $L\times L$ response matrix $A^{(r)}$, adapted to the wider energy bins of the multi-comparator system with a limited number of energy thresholds. The inversion of the reduced response matrix, or the iterative solution of Eq.16, allows to retrieve the number of input events $m_{i}$.
@@ 2.3 Simulation of a CdTe detector
In order to validate the method described in the previous sections, the response of a pixelated CdTe detector of 1 mm thickness and segmented in 9x9 pixels was simulated with the Geant4 toolkit (Agostinelli et al., 2003) version 11.1.3. Different pixel sizes were considered, between 100 and 300 µm. X-ray photons were generated with a direction perpendicular to the detector, and with an energy randomly selected from a polychromatic X-ray spectrum produced with the Spektr code (Punnoose et al., 2016). The photon source was placed at a fixed distance from the detector with transversal coordinates uniformly distributed to illuminate the whole detector.
@@ 3.1 Simulation studies
Figure 5: MAPE values from the comparison of the CBRM correction (left) and the ACS correction (right) with the reference spectrum of an ideal detector, as a function of the pixel size and the noise discrimination threshold. The input spectrum from a 120 kV X-ray tube was attenuated by 120 mm of H2O and 0.2 mm of Iodine, and a random noise of 1 keV rms was added to the output of each pixel. The results were obtained from Geant4 simulations.
The effectiveness of the CBRM correction depends also on the number of energy bins used for the definition of the response matrix. In general, a high number of energy bins is desirable for an accurate determination of the matrix; this can be obtained by changing the thresholds of the two discriminators of Fig.1(right) in small steps. However, the typical number of comparators and the number of energy bins available in standard pulse-height acquisition systems is limited. The procedure described in Section2.2.3 can be followed to reduce the response matrix to match the number of energy bins of a standard acquisition.
@@
3.2 Study of experimental data
The analysis of the data collected at the SYRMEP beamline with the silicon detector and the Timepix4 ASIC was performed using energy bins of 0.5 keV width. The response matrix was determined by merging all the data collected at 18 different energies, to define the splitting probabilities over a wide energy range. The CBRM matrix was therefore applied to recover the spectrum of each single acquisition, and the results were compared with the spectra obtained from a clustering algorithm based on 3x3 pixel blocks (Sect.2.5).